# Optical aberrations in autonomous driving: Physics-informed parameterized temperature scaling for neural network uncertainty calibration

### v1

@@ 5 Experiments and results
5.1 Multi-task network performance and the insufficiency of the half-Nyquist criterion
The performance w.r.t. the target application, semantic segmentation of the A2D2 dataset, is quantitatively evaluated in terms of the mIoU. The peak mIoU on the test dataset amounts to . The segmentation performance is qualitatively visualized by a random instance in Figure 6.
Furthermore, Figure 5 depicts the mIoU and the mECE as a function of different optical metrics. It is evident that the peak performance is not given in the absence of optical aberrations (diffraction-limited case), which might be counterintuitive at first glance. The mIoU is maximized for instances that reflect the mean-level of optical aberrations within the training dataset. As a consequence, it is of paramount importance to incorporate the optical aberrations within the perception chain proportionally to their occurrence in part-level measurements. By doing so, the augmented training dataset will be centered at the expected optical quality of the produced perception chain, and the mIoU as well as the mECE will be implicitly tuned for this aberration scenario.
This conclusion aligns well with previous work in the field of deep optics (Chang and Wetzstein,, 2019; Tseng et al.,, 2021; Yang et al.,, 2023), where the perception chain is holistically optimized alongside the neural network training. This is done by constructing a differentiable, physics-based optics model and by assigning the corresponding optical parameters as trainable hyperparameters. The contributions in this field (Chang and Wetzstein,, 2019; Tseng et al.,, 2021; Yang et al.,, 2023) strongly indicate that optical quality is not all what you need. To the best of our knowledge, this result has only been shown with respect to the target application performance, e.g., image classification, depth estimation, 3D object detection etc. With our work, we demonstrate that this effect also holds true in terms of the neural network calibration performance.
@@ 5.1 Multi-task network performance and the insufficiency of the half-Nyquist criterion
The regression function is parameterized by the coefficient vector and the combined uncertainty will be determined by applying the multivariate law of uncertainty propagation (Ludwig,, 2023):
The covariance matrix for the parameters is calculated by a Monte-Carlo study (Raychaudhuri,, 2008) considering the batch-wise standard deviation of the mIoU and the mECE respectively. In total regression curves were calculated, wherefore the extension factor is given by for a confidence level of (for Guides in Metrology, JCGM; Pesch,, 2003). Finally, the symmetrical interval spanned by the combined uncertainty determines the confidence bands around the regression function and is illustrated in Figure 5 in gray.
The explanatory power of the postulated regression function (10) is further quantified in terms of the unexplainable variance, which is given by the ratio of the mean squared error (MSE) (Upton and Cook,, 2008) over the variance of the dataset itself. Statistically, the MSE measures the ensemble spread around the regression line. Benchmarking regression models according to the unexplainable variance is favorable because it effectively measures how well the regression model outperforms the naive estimate given by the arithmetic mean (Wolf et al., 2023a, ). In addition, if the complexity of the regression model is increased the degrees of freedom in the MSE computation decreases, which acts as a penalty for more complex regression models like in Ridge regression (Taboga,, 2021; Theobald,, 1974; Farebrother,, 1976). The quantitative values for the unexplainable variance are presented in Figure 5 and indicate that the model function does not significantly outperform the arithmetic mean in case of the MTF. On the other hand, the model function is superior for the Strehl ratio and the OIG. Consequently, the coefficient vector can be interpreted as a sensitivity vector in case of the Strehl ratio and the OIG. Of special importance are for the linear performance drift in the low-aberration regime and the coefficients and for the exponential decay in the large-aberration regime. These sensitivity coefficients can be employed for determining valid operational domains for the perception chain and for comparing the robustness across different neural network architectures.
@@
Nevertheless, it is evident, that the MTF at half Nyquist frequency fits the regression function well in the low-aberration regime but the huge spread in the codomain for the large-aberration regime leads to a wide confidence band around the global trend. The correlation of the MTF breaks down for severe optical aberrations because of the monotonicity violation in the large-aberration regime. As long as the MTF function is monotonically decreasing, the value of the MTF function at half Nyquist frequency will correlate with the area enclosed by the MTF curve. Since the area under the MTF curve, which equals the Strehl ratio if normalized by the diffraction-limited case, shows a robust correlation to the mIoU across the entire spatial frequency domain, the MTF at half Nyquist frequency is a valid optical performance indicator as long as the MTF curve is monotonically decreasing. As the optical aberrations in the automotive industry are typically too severe in magnitude to satisfy the monotonicity constraint, the MTF at half Nyquist frequency is not a suitable measure for safeguarding AI-based autonomous driving algorithms against optical perturbations. Consequently, system MTF requirements in the large-aberration regime should be considered as invalid.
5.2 PIPTS calibration quality
@@
The calibration performance is visualized by the reliability diagrams in Figure 7. The multi-task network is inherently well calibrated, wherefore TS results in an identity mapping for the confidences with . The PTS and PIPTS instance-wise calibrators outperform TS significantly (see below) by over . This highlights the gain in expressive power provided by the superior information capacity of the post-hoc CNN-based calibrators (Tomani et al.,, 2022). The physical inductive bias of PIPTS results in an additional mECE boost of roughly for the diffraction-limited case. If aberrations are considered, then the performance boost increases, as can be seen by the reduced spread of the red triangles especially in the low-confidence bins.
A valid question might be raised about how significant this performance boost is. In order to tackle this question, we utilized the Deep ensemble approach (Lakshminarayanan et al.,, 2016). An ensemble of 11 PIPTS models has been trained with the same hyperparameters (congruent loss function landscape in the parameter space) but random and hence different weights initialization. By doing so, the ensemble mean is determined and its corresponding standard deviation is utilized as an predictive uncertainty estimator (Lakshminarayanan et al.,, 2016). The significance level of the PIPTS performance gain has been set to utilizing the Student t-distribution and a corresponding extension factor of for an ensemble with degrees of freedom (for Guides in Metrology, JCGM; Pesch,, 2003). For each ensemble member, the robustness analysis presented in Figure 5 is repeated and the corresponding global trend for the well-correlating optical metrics, the Strehl ratio and the OIG, is extracted by quantifying the regression function parameters . Subsequently, the mean performance boost and the corresponding standard deviation for the mean are evaluated and visualized.
@@
Figure 8 depicts the mECE curves for the PTS and PIPTS calibrator against the TS calibration results for the Strehl ratio. Those curves do not represent functional relationships because the calibration performance for all three post-hoc techniques is maximized for the mean-aberration regime. The PIPTS performance boost is not significant for the diffraction-limited case if a confidence level of is considered in the view of the Student t-distribution. The performance boost increases almost linearly with the aberration magnitude until the mean-aberration regime is approached. Within the dataset centroid, the multi-task model seems to be inherently better calibrated, wherefore the supplementary information about the Zernike coefficients is not as beneficial as for long-tail samples with fewer aberrations in terms of the Strehl ratio. This conclusion also underpins the observation, that the performance boost of PIPTS increases again after passing the mode value of the augmented dataset distribution in the large-aberration regime. A similar outcome is observed if the OIG is considered as the optical target metric, as illustrated in Figure 9. As a result, the PIPTS performance boost is significant on a confidence level for the mean- and large-aberration regime but not for the low-aberration regime close to the diffraction limit, as the physical prior does not add significant information within this domain.
@@
The performance boost of PIPTS – generated by the physical inductive bias – manifests itself as a slight distribution shift in the predicted temperature over the entire test dataset. As a toy-example, Figure 10 depicts the histograms of the predicted temperature deviation for the aberration scenario: , , . The temperature deviation is given as the difference between the instance-wise temperature predicted by the corresponding calibrator (PTS or PIPTS) and the instance-wise optimal temperature, which is obtained by minimizing the mECE as a function of the temperature for each instance separately. The distributions show a non-zero bias, which is subject to the dataset shift induced by the optical aberrations. The bias is significantly reduced if the PIPTS calibrator is employed instead of the PTS calibrator. The parameters of the underlying Gaussian distribution (bias and standard deviation ) were determined by a negative log-likelihood fit and the corresponding parameter uncertainty is quantified by the local curvature of the negative log-likelihood curve according to the Fisher information.
6 Benefits for autonomous driving
@@
The predicted, effective wavefront aberrations of the overall system can not only be used as an inductive bias for the PIPTS calibrator but also for end-of-line testing. The multi-task network enables the end-of-line testing by absorbing the non-linear interplay of the optical aberrations induced by the ADAS camera and the windshield into the information capacity of the multi-task model. The non-linear mapping of the part-level Zernike coefficient vectors of the ADAS camera and the windshield to the Zernike coefficient vector of the entire perception chain permits individual part testing at the supplier site, as it is essential for automotive industry processes according to the Vee-model (on Systems Engineering, INCOSE). The predicted Zernike coefficient vector of the overall optical system can be measured end-of-line (Wolf et al., 2024b, ; Wolf et al., 2023a, ) and serves as the ground truth for the regression task.
7 Conclusion
@@
Our contribution manifests itself threefoldly. First, we quantitatively demonstrated that the Strehl ratio and the OIG outperform the MTF at half Nyquist frequency in terms of the Chatterjee rank correlation measure. Secondly, we showed experimental evidence on the superiority of PIPTS over PTS, which implies that incorporating physical priors to the PTS calibrator enhances its expressive power. Finally, we highlighted the benefits of the coupled decoder head for the Zernike coefficient prediction in the light of system-level requirements. The multi-task network provides a tool for capturing the non-linear mapping from the Zernike coefficient vectors of the ADAS camera and the windshield to the system-level wavefront aberration map. This enables the automotive industry to derive part-specific requirements as the verification strategy according to the Vee-model is demanding. In a nutshell, our contribution paves the way for establishing trustworthiness and robustness in AI-based autonomous driving functionalities by ensuring superior confidence calibration under optical aberrations in the perception chain and by providing a physical sound toolchain for deriving part-specific optical requirements.
Data availability statement: The A2D2 semantic segmentation dataset from AUDI, used for training and evaluation, is publicly available under the CC BY-ND 4.0 license on https://www.a2d2.audi. Our code is not publicly available in order to protect industrial property rights.

### v2

@@ V Experiments and results
V-A Multi-task network performance and the insufficiency of the half-Nyquist criterion
The performance w.r.t. the target application, semantic segmentation of the A2D2 dataset, is quantitatively evaluated in terms of the mIoU. Across the entire test dataset, the peak average mIoU evaluates to . Furthermore, the segmentation performance is qualitatively visualized by a random instance in Figure 6.
Moreover, Figure 5 depicts the mIoU and the mECE as a function of different optical metrics. It is evident that the peak performance is not given in the absence of optical aberrations (diffraction-limited case), which might be counterintuitive at first glance. The mIoU is maximized for instances that reflect the mean-level of optical aberrations within the training dataset. As a consequence, it is of paramount importance to incorporate the optical aberrations within the perception chain proportionally to their occurrence in part-level measurements. By doing so, the augmented training dataset will be centered at the expected optical quality of the produced perception chain, and the mIoU as well as the mECE will be implicitly tuned for this aberration scenario.
This conclusion aligns well with previous work in the field of deep optics [61, 62, 63], where the perception chain is holistically optimized alongside the neural network training. This is done by constructing a differentiable, physics-based optics model and by assigning the corresponding optical parameters as trainable hyperparameters. The contributions in this field [61, 62, 63] strongly indicate that optical quality is not all what you need. To the best of our knowledge, this result has only been shown with respect to the target application performance, e.g., image classification, depth estimation, 3D object detection etc. With our work, we demonstrate that this effect also holds true in terms of the neural network calibration performance.
@@ V-A Multi-task network performance and the insufficiency of the half-Nyquist criterion
The regression function is parameterized by the coefficient vector and the combined uncertainty will be determined by applying the multivariate law of uncertainty propagation [64]:
The covariance matrix for the parameters is calculated by a Monte-Carlo study [65] considering the batch-wise standard deviation of the mIoU and the mECE respectively. In total regression curves were calculated, wherefore the extension factor is given by for a confidence level of [66, 67]. Finally, the symmetrical interval spanned by the combined uncertainty determines the confidence bands around the regression function and is illustrated in Figure 5 in gray.
The explanatory power of the postulated regression function (10) is further quantified in terms of the unexplainable variance, which is given by the ratio of the mean squared error (MSE) [21] over the variance of the dataset itself. Statistically, the MSE measures the ensemble spread around the regression line. Benchmarking regression models according to the unexplainable variance is favorable because it effectively measures how well the regression model outperforms the naive estimate given by the arithmetic mean [47]. In addition, if the complexity of the regression model is increased the degrees of freedom in the MSE computation decreases, which acts as a penalty for more complex regression models like in Ridge regression [68, 69, 70]. The quantitative values for the unexplainable variance are presented in Figure 5 and indicate that the model function does not significantly outperform the arithmetic mean in case of the MTF. On the other hand, the model function is superior for the Strehl ratio and the OIG. Consequently, the coefficient vector can be interpreted as a sensitivity vector in case of the Strehl ratio and the OIG. Of special importance are for the linear performance drift in the low-aberration regime and the coefficients and for the exponential decay in the large-aberration regime. These sensitivity coefficients can be employed for determining valid operational domains for the perception chain and for comparing the robustness across different neural network architectures.
@@
Nevertheless, it is evident, that the MTF at half Nyquist frequency fits the regression function well in the low-aberration regime but the huge spread in the codomain for the large-aberration regime leads to a wide confidence band around the global trend. The correlation of the MTF breaks down for severe optical aberrations because of the monotonicity violation in the large-aberration regime. As long as the MTF function is monotonically decreasing, the value of the MTF function at half Nyquist frequency will correlate with the area enclosed by the MTF curve. Since the area under the MTF curve, which equals the Strehl ratio if normalized by the diffraction-limited case, shows a robust correlation to the mIoU across the entire spatial frequency domain, the MTF at half Nyquist frequency is a valid optical performance indicator as long as the MTF curve is monotonically decreasing. As the optical aberrations in the automotive industry are typically too severe in magnitude to satisfy the monotonicity constraint, the MTF at half Nyquist frequency is not a suitable measure for safeguarding AI-based autonomous driving algorithms against optical perturbations. Consequently, system MTF requirements evaluated at half-Nyquist frequency should be considered as invalid for the large-aberration regime.
V-B PIPTS calibration quality
@@
The calibration performance is visualized by the reliability diagram in Figure 7 for the baseline model calibrated with TS. As the multi-task network is inherently well calibrated for the diffraction-limited case by utilizing class-balancing and loss-focusing during training, the optimal temperature according to TS is given by , which corresponds to an identity mapping of the confidences. In addition, Figure 7 visualizes the reliability diagrams for the multi-task network after applying PTS and PIPTS, respectively. It is evident that the instance-wise calibrators significantly outperform TS by over . This highlights the gain in expressive power provided by the superior information capacity of the post-hoc CNN-based calibrators [11]. The physical inductive bias of PIPTS results in an additional mECE boost of roughly for the diffraction-limited case. If aberrations are considered, then the knowledge about the dataset shift magnitude successfully counteracts the calibration bias. In addition to accuracy improvements, PIPTS also slightly improves the calibration precision, which is notable by narrower bin-wise calibration error distributions.
A valid question might be raised about how significant this performance boost is. In order to tackle this question, we utilized the Deep Ensemble approach [23]. An ensemble of 11 PIPTS models has been trained with the same hyperparameters (congruent loss function landscape in the parameter space) but random and hence different weights initialization. Each ensemble member is used to calibrate the baseline model and the robustness analysis presented in Figure 5 is repeated for the well-correlating optical metrics, the Strehl ratio and the OIG. Consequently, the violation of the calibration condition [10] under optical aberrations can be quantified by extracting the regression function parameters . Subsequently, the mean performance boost across all ensemble members and the corresponding standard deviation for the mean are evaluated and visualized in Figure 8(a) and 8(b). Utilizing the standard deviation of the mean calibration performance in terms of the mECE as a predictive uncertainty estimator for the PIPTS calibrator implicitly neglects the aleatoric uncertainty component. If it is intended to include this component into the uncertainty estimation regarding the mECE, second order distributions have to be considered. This could be done by extending the output space of the PIPTS calibration network such that the conditional uncertainty of the temperature is predicted alongside the instance-wise calibration temperature. The conditional uncertainty can then be propagated to retrieve an uncertainty estimate for the mECE. Unfortunately, this approach would require us to adjust the loss function. If a Gaussian distribution is assumed to characterize the noise in the training dataset of the PIPTS calibrator, then the Gaussian negative log-likelihood loss [71] could be employed for training. The advantage of learning the conditional uncertainty alongside the instance-wise calibration temperature also comes with a downside, as the Gaussian negative log-likelihood loss relies on ground truth temperatures. As this information is not available a priori, the optimal temperature would need to be retrieved before training for every instance by minimizing the mECE. In order to avoid a supervised regression setup for the PIPTS calibration network training, we employ the modulated mECE directly as a loss function, which eliminates the need of ground truth temperatures. As a consequence, we will restrict our study to first order distributions and the PIPTS calibrator exclusively predicts the temperature. Neglecting the aleatoric uncertainty component for the evaluation of the PIPTS calibration performance is reasonable, as the aleatoric uncertainty component essentially measures the quality of the baseline segmentation model because of the noise that the baseline model generates in its predictions due to misclassifications. In order to avoid this dependency on the performance of the segmentation model, only the epistemic uncertainty estimate is used for assessing the predictive uncertainty of the PIPTS calibration network. The significance level of the PIPTS performance gain has been set to utilizing the Student’s t-distribution and a corresponding extension factor of for an ensemble with degrees of freedom [66, 67]. Here, the Student’s t-distribution has to be considered, as the standard deviation is estimated based on a very limited number of degrees of freedom [66, 67].
@@
Figure 8(a) depicts the mECE curves for the PTS and PIPTS calibrator against the TS calibration results for the Strehl ratio. Those curves do not represent functional relationships because the calibration performance for all three post-hoc techniques is maximized for the mean-aberration regime. The PIPTS performance boost is not significant for the diffraction-limited case if a confidence level of is considered in the view of the Student’s t-distribution. The performance boost increases almost linearly with the aberration magnitude until the mean-aberration regime is approached. As the optical degradations approach the level of the mean optical quality of the training dataset, given by the dataset centroid, the benefit of employing PIPTS decreases. Hence, the supplementary information about the Zernike coefficients is not as beneficial as for long-tail samples with fewer aberrations in terms of the Strehl ratio. This might be the case if a neural network trained on augmented data is used to infer information from synthetic data, which typically does not comprise optical aberrations of the windshield. On the other hand, in the large-aberration regime, the performance boost of PIPTS increases again after passing the mode value of the augmented dataset distribution. This scenario might occur if aging effects of the windshield are considered.
A similar outcome is observed if the OIG is considered as the optical target metric, as illustrated in Figure 8(b). As a result, the PIPTS performance boost is significant on a confidence level for the mean- and large-aberration regime but not for the low-aberration regime close to the diffraction limit, as the physical prior does not add significant information within this domain.
@@
The performance boost of PIPTS – generated by the physical inductive bias – manifests itself as a slight distribution shift in the predicted temperature over the entire test dataset. As a toy-example, Figure 9 depicts the histograms of the predicted temperature deviation for the aberration scenario: , , and . The temperature deviation is given as the difference between the instance-wise temperature predicted by the corresponding calibrator (PTS or PIPTS) and the instance-wise optimal temperature, which is obtained by minimizing the mECE as a function of the temperature for each instance separately. The distributions show a non-zero bias, which is subject to the dataset shift induced by the optical aberrations. The bias is significantly reduced if the PIPTS calibrator is employed instead of the PTS calibrator. The parameters of the underlying Gaussian distribution (bias and standard deviation ) were determined by a negative log-likelihood fit and the corresponding parameter uncertainty is estimated by the inverse of the local curvature of the negative log-likelihood curve according to the Cramer-Rao-Frechet bound [72, 73, 74].
VI Benefits for autonomous driving
@@
The predicted, effective wavefront aberrations of the overall system can not only be used as an inductive bias for the PIPTS calibrator but also for end-of-line testing. The multi-task network enables the end-of-line testing by absorbing the non-linear interplay of the optical aberrations induced by the ADAS camera and the windshield into the information capacity of the multi-task model. The non-linear mapping of the part-level Zernike coefficient vectors of the ADAS camera and the windshield to the Zernike coefficient vector of the entire perception chain permits individual part testing at the supplier site, as it is essential for automotive industry processes [75] according to the Vee-model [76].
VII Conclusion
@@
Our contribution manifests itself threefoldly. First, we quantitatively demonstrated that the Strehl ratio and the OIG outperform the MTF at half Nyquist frequency in terms of the Chatterjee rank correlation measure. Secondly, we showed experimental evidence on the superiority of PIPTS over PTS, which implies that incorporating physical priors to the PTS calibrator enhances its expressive power. Finally, we highlighted the benefits of the coupled decoder head for the Zernike coefficient prediction in the light of system-level requirements. The multi-task network provides a tool for capturing the non-linear mapping from the Zernike coefficient vectors of the ADAS camera and the windshield to the system-level wavefront aberration map. This enables the automotive industry to derive part-specific requirements as the verification strategy according to the Vee-model is demanding [76, 77]. In a nutshell, our contribution paves the way for establishing trustworthiness and robustness in AI-based autonomous driving functionalities by ensuring superior confidence calibration under optical aberrations in the perception chain and by providing a physical sound toolchain for deriving part-specific optical requirements.
Dominik Werner Wolf studied physics at the Karlsruhe Institute of Technology and at the Ruprecht-Karls-University of Heidelberg. His academic journey led him to CERN, where he conducted research on tune modulations at the LHC for his Bachelor’s thesis and worked on novel beam stabilization techniques for his Master’s thesis. At the OPMD lab of the University of Oxford, he contributed to the Mu3e detector development and the assembly strategy, showcasing interdisciplinary prowess. With over six years of industry experience in metrology and neural network robustness, he bridges theory and application seamlessly. His passion lies in computational physics, tackling problems across disciplines with relentless curiosity. |
Alexander Braun has 20 years of experience with optical technologies of all flavours, from fundamental quantum optics research to optical quality questions arising during assembly of 100.000 cameras per year. During his position as optical designer at Kostal (Automotive Tier 1) he was responsible for the optical quality of the camera ADAS, designing MTF test benches for all purposes. As a Professor for Physics at the University of Applied Sciences in Duesseldorf he is now looking at all the questions surrounding optical quality from a scientific point of view. Special research interests are the limits of current image quality algorithms and linking optical design to function performance. |
Markus Ulrich is a distinguished academic and industry expert with over twenty years of experience bridging the realms of academia and industry. Currently a Professor for Machine Vision Metrology at the Institute of Photogrammetry and Remote Sensing at the Karlsruhe Institute of Technology (KIT), his journey began with a PhD degree at the Department of Civil, Geo, and Environmental Engineering of Technical University of Munich (TUM). Prior to academia, he served as the Invention and Patent Manager at MVTec Software GmbH, Munich, and headed the research team at the same institution. In 2017, he completed his habilitation (venia legendi) and was appointed as a Privatdozent (Lecturer) at KIT. His research interests include machine vision, close-range photogrammetry, machine learning, and their applications in industry for quality inspection, automation, and robotics. |